Sarah decided to try telemarketing. She calls people from 3pm to 9pm. Marley has a 15% chance of being called during this time. If Marley eats dinner from 6:6:30, how likely is it that she'll have a conversation with Sarah during this time?
Answer by jrfrunner(365)

let A=Marley is called during 3pm to 9pm
let B=Sarah calls someone during 6pm to 6:30pm
--
Given P(A)=0.15
---
P(B)=30/360 =1/12 since there are 30 minutes during the 6hours that he can be called.
-
Need to know P(AnB) if we assume A & B are independent
then P(AnB)=P(A)*P(B)=0.15*1/12=0.0125